HUBBLE SPACE TELESCOPE DAILY REPORT #2151 PERIOD COVERED: 0000Z (UTC) 06/23/98 - 0000Z (UTC) 06/24/98 Daily Status Report as of 175/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7928 (STIS Spectroscopic CCD Flatfielding -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat field calibration images in the STIS spectroscopic mode. The calibrations were completed as planned, and no problems were reported. 1.2 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Two Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. The observations were completed as scheduled, and no anomalies were reported. 1.7 Completed Three Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC/3 7843 (Testing Protostellar Collapse Theory through Extinction Mapping) The Near Infrared Camera (Camera Three) was used to make observations for extinction mapping of the protostellar star B335. The identification of collapsing protostars remains a "holy grail" of star formation studies. The best collapse candidate is generally recognized to be B335, an isolated round globule containing a deeply embedded low luminosity young stellar object. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Operations Notes: The SSR EDAC error counter was cleared two times per ROP SR-1. Operations Request 13840 directed the start of the Battery 4 Capacity Test at 174/1702Z. Discharge began at 174/2001Z. A real-time map was commanded at 175/0520Z. Using ROP IC-2, the SSA transmitter was turned off or on during two intervals: 175/0554Z-0607Z and 175/0650Z-0709Z. Per Operations Note 446, an extra DF-224 was requested and cpatured on CCS strings A and B at 175/0653Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH